# A Simple and Effective $L_{2}$ Norm-Based Strategy for KV Cache Compression

### v2

@@ 4 Experiments
We use FastGen (Holmes et al., 2024), a popular method for KV Cache compression, as a baseline for assessing the effectiveness of our method. It is important to note that, like the majority of methods in the literature, FastGen utilises attention scores, which makes it incompatible with the popular FlashAttention (Dao et al., 2022), thereby limiting its efficiency and usability. For a fair comparison, we implement FastGen without using the attention scores, i.e., we only consider local, punctuation and special tokens. We perform experiments on language modelling with the Llama3 model (Dubey et al., 2024). Our method still outperforms FastGen with up to 50% KV Cache eviction. We show the results in Figure 6.
@@ 5 Analysis
Interestingly, this sparsity aligns with the concept of "sink" tokens, as identified in previous studies (Xiao et al., 2024). These tokens capture a direction in the embedding space such that many queries align closely with it, leading to increased attention scores for these tokens. In simpler terms, when the key embeddings of certain tokens are dominated by a small number of dimensions, it appears to create a situation where a large number of queries - regardless of their specific content - are naturally drawn to these tokens, increasing their attention weight.
@@
We hypothesise that the lower $L_{2}$ norm reflects a partial use of the available embedding space, leading to increased attention for these tokens. To test this, we zeroed out the dimensions responsible for the peaked activations in low-norm key embeddings and observed significant changes in attention maps (Figure 9). Randomly altering dimensions did not produce the same effect, suggesting that the specific active dimensions play a critical role in attention distribution. This finding suggests that the $L_{2}$ norm may serve as a proxy for the extent to which an embedding utilises the available vector space and, consequently, the degree to which it influences attention. Lower $L_{2}$ norm appears to correspond to embeddings that drive disproportionately high attention values due to their alignment with a common "sink" direction.
@@ 6 Related Work
Recently, various long-context LLMs, such as Gemini-Pro-1.5 (Reid et al., 2024), Claude-3 (Anthropic, 2024), and GPT4 (Achiam et al., 2023), have shown the promising capability to process hundred thousands of tokens in the context. The increased number of input lengths results in a high decoding latency; thus, there has been a growing interest in speeding up the decoding with long contexts. Some works propose efficient memory management strategies to reduce the IO time overheads, e.g., PageAttention (Kwon et al., 2023), Infinite-LLM (Lin et al., 2024) and vAttention (Prabhu et al., 2024). Another line of research focuses on compressing the KV Cache to improve efficiency. DMC (Nawrot et al., 2024) compresses KV Cache by dynamically merging tokens while requiring expensive continual pre-training. For fine-tuning free compression strategy, H2O (Zhang et al., 2024b) identifies important KV pairs by leveraging the attention scores from all queries, FastGen (Ge et al., 2023a) leverages the different attention patterns in different heads for compression, and SnapKV (Li et al., 2024) selects KV pairs based on attention scores from user’s query. Unlike these works, our method only utilises the $L_{2}$ norm of embedding for compression without leveraging the attention information, and to the best of our knowledge, we are the first to find that the influence of a KV pair can be determined by $L_{2}$ norm. Previous work (Darcet et al., 2024) finds the hidden states with high $L_{2}$ norm usually aggregate more important and global information. On the other hand, our findings indicate that a low $L_{2}$ norm of key embedding generally results in a high attention score.

### v4

@@ 4 Experiments
We use FastGen (Ge et al., 2023a), a popular method for KV Cache compression, as a baseline for assessing the effectiveness of our method. It is important to note that, like the majority of methods in the literature, FastGen utilises attention scores, which makes it incompatible with the popular FlashAttention (Dao et al., 2022), thereby limiting its efficiency and usability. For a fair comparison, we implement FastGen without using the attention scores, i.e., we only consider local, punctuation and special tokens. We perform experiments on language modelling with the Llama3 model (Dubey et al., 2024). Our method still outperforms FastGen with up to 50% KV Cache eviction. We show the results in Figure 6.
@@ 5 Analysis
Interestingly, this sparsity aligns with the concept of "sink" tokens, as identified in previous studies (Xiao et al., 2024). These tokens capture a direction in the embedding space such that many queries align closely with it, leading to increased attention scores for these tokens. Specifically, when the key embeddings of certain tokens are dominated by a limited set of dimensions, they create a focal point, attracting a wide range of queries – regardless of their individual content – and amplifying their attention weights.
@@
We hypothesise that the lower $L_{2}$ norm reflects a partial use of the available embedding space, leading to increased attention for these tokens. To examine this hypothesis, we zeroed out the dimensions responsible for the peaked activations in low-norm key embeddings and observed significant changes in attention maps (Figure 9). In contrast, altering random dimensions did not produce the same effect, highlighting the importance of these specific dimensions. This finding suggests that the $L_{2}$ norm may serve as a proxy for the extent to which an embedding utilises the available vector space and, consequently, the degree to which it influences attention. Lower $L_{2}$ norm appears to correspond to embeddings that drive disproportionately high attention values due to their alignment with a common direction. Cancedda (2024) offers additional insight into this phenomenon, suggesting that attention sinks engage with other tokens through a “dark” subspace within the embedding space.
@@ 6 Related Work
Recently, various long-context LLMs, such as Gemini-Pro-1.5 (Reid et al., 2024), Claude-3 (Anthropic, 2024), and GPT4 (Achiam et al., 2023), have shown the promising capability to process hundred thousands of tokens in the context. The increased number of input lengths results in a high decoding latency; thus, there has been a growing interest in speeding up the decoding with long contexts. Some works propose efficient memory management strategies to reduce the IO time overheads, e.g., PageAttention (Kwon et al., 2023), Infinite-LLM (Lin et al., 2024) and vAttention (Prabhu et al., 2024). Another line of research focuses on compressing the KV Cache to improve efficiency. DMC (Nawrot et al., 2024) compresses KV Cache by dynamically merging tokens while requiring expensive continual pre-training. For fine-tuning free compression strategy, H2O (Zhang et al., 2024b) identifies important KV pairs by leveraging the attention scores from all queries, FastGen (Ge et al., 2023a) leverages the different attention patterns in different heads for compression, and SnapKV (Li et al., 2024) selects KV pairs based on attention scores from user’s query. Unlike these works, our method only utilises the $L_{2}$ norm of embedding for compression without leveraging the attention information, and to the best of our knowledge, we are the first to find that the influence of a KV pair can be determined by $L_{2}$ norm. Previous work (Darcet et al., 2024) finds the hidden states with high $L_{2}$ norm usually aggregate more important and global information. On the other hand, our findings indicate that a low $L_{2}$ norm of key embedding generally results in a high attention score. Concurrently to this work, Guo et al. (2024) uses the $L_{1}$ norm of values in the KV Cache and attention scores for compression.
@@ 8 Limitations
9 Acknowledgments
Alessio Devoto was supported by Sapienza Grant RM1221816BD028D6 (DeSMOS). Yu Zhao was partly supported by the UKRI Centre for Doctoral Training in Natural Language Processing, funded by UK Research and Innovation (grant EP/S022481/1) and the University of Edinburgh, School of Informatics. Pasquale Minervini was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.